Write the value of tan 75°.

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
tan 75=tan(45+30)
=\frac{tan 45+tan 30}{1-tan45 tan 30}
=\frac{1+1/\sqrt{3}}{1- 1/\sqrt{3}}
=\frac{\sqrt{3}+1}{\sqrt{3}-1}
Thanks and regards
Shaik Aasif
askIITians faculty